# Dark matter cooling during early matter-domination boosts sub-earth halos

### v1

@@ 4.2 Model II: p-wave scattered fermion dark matter
Figure 8: The allowed parameter region satisfying $m_{\chi}<M\leq 2m_{\chi}$ and reproducing the present day relic density of dark matter for a fixed value of $T_{\rm RH}=100$ MeV. In the dark shaded region, dark matter kinetically decouples during the EMDE, while it decouples during radiation dominated epoch in the light-shaded region. Enhancement in sub-earth halo population can not be observed in the region on the left side of the cyan dashed curve, even though DM remains fully kinetically decoupled during EMDE. The isocontours of the free-streaming horizon are shown by white lines.
@@
  • The DM shows full kinetic decoupling during the EMDE in the dark coloured area, while the relic is staisfied in the entire parameter space within the specified boundaries.
  • The isocontours of free-streaming horizon are projected as white curves on the allowed parameter space. Due to the full decoupling of the DM, the value of $\lambda_{\rm fsh}$ is approximately two orders of magnitude smaller compared to a similarly situated benchmark point in the Fig. 7, where the DM decouples only partially.
  • The region left to the cyan dot-dashed line does not exhibit boost in the population of sub-earth halos even if DM becomes fully decoupled during the EMDE. This is because the free-streaming horizon of the DM in that region exceeds the cut-off for the corresponding $T_{\rm RH}$.
@@
We illustrate the differential halo mass function at several redshift values for both the Model I and II in Fig. 9, keeping $T_{\rm RH}=10$ MeV and $m_{\chi}=95$ GeV for both cases. The relevant couplings are chosen, as in Fig. 3, to satisfy the present day relic density. For this benchmark point, the DM decouples around $T=2.2$ GeV for both Models. However, the free-streaming horizon of the DM is $\lambda_{\rm fsh}=0.264$ pc for Model I and $0.019$ pc for Model II, respectively. The smaller $\lambda_{\rm fsh}$ in Model II is attributed to the extra cooling due to full kinetic decoupling before the end of reheating. The additional cooling in Model II is further confirmed by the DM temperature at $a_{\rm RH}$, with $T_{\chi}(a_{\rm RH})$ being $55.5$ KeV for Model I and $12$ eV for Model II. As shown in Fig. 9, the differential halo mass function peaks around $M=0.1\,M_{\oplus}$ for both models, while the absolute value is higher in Model II due to the reduced $\lambda_{\rm fsh}$.
Figure 9: Differential halo mass fraction is demonstrated for both the models at different redshifts values. For the benchmark points as mentioned in the caption of Fig. 3, the free-streaming horizon of DM turns out to be $0.264$ $pc$ and $0.019$ $pc$ for Model I and Model II, respectively. The differential halo mass function peaks around $M=0.1\,M_{\oplus}$.
@@ 5 Conclusions
We have comprehensively studied the chemical and kinetic decoupling of thermal dark matter in the presence of a pre-BBN early matter-dominated era. Specifically, we explore a scenario where a long-lived scalar field dominated the energy density of the universe, eventually transferring energy to the radiation through a temperature-dependent dissipation rate ($\Gamma_{\phi}\sim T$). We demonstrate that the kinetic decoupling of the DM during the EMDE is partial if the elastic scattering of the DM with the bath particles is dominated by $s$-wave partial wave amplitudes, while the DM fully decouples during the EMDE if it undergoes $p$-wave scattering. This is in contrast to an EMDE with constant rate of energy dissipation ($\Gamma_{\phi}\sim{\rm const.}$), where $s$-wave scatterings are inefficient to kinetically decouple the DM, while $p$-wave scatterings can partially decouple it from the thermal bath.
@@
An early kinetic decoupling reduces the free-streaming horizon of the DM ($\lambda_{\rm fsh}$) due to additional cooling of the DM during the EMDE. In addition, the presence of an EMDE also enhances the matter perturbations for scales entering the horizon during this epoch. The population of the DM halos with masses smaller than $M_{\rm RH}$ receive a boost due to the combined effect of reduced free-streaming horizon and enhanced matter perturbations. The boost in the number of small scale DM halos may in turn lead to an increase in the gamma-ray annihilation signals from the DM halos, offering an interesting probe for the early matter dominated cosmology.
@@
However, kinetic decoupling of the DM during the EMDE does not guarantee an enhanced structure growth. If the annihilation of the DM which determines its relic and the elastic scattering with bath particles are governed by the same interaction strength, the free-streaming horizon reduces in the EMDE. On the other hand, in presence of multiple interactions which decorrelate the mechanisms for chemical and kinetic decoupling, free-streaming horizon may not be reduced in presence of the EMDE, unless the kinetic decoupling temperature is greater than a threshold value compared to the reheating temperature, see Fig. 1.
@@
Moreover, we have shown that the formation of DM halos receive a boost iff the free-streaming horizon is smaller than a certain cut-off ($\lambda^{c}_{\rm fsh}$). We have quantified $\lambda^{c}_{\rm fsh}$, which depends on the reheating temperature, by comparing $\sigma(M)$ in the presence of EMDE and in the purely radiation dominated universe with $\lambda_{\rm fsh}=0$.
@@
As mentioned above, if the chemical and kinetic decoupling are correlated by the same interaction strengths, the requirement of kinetic decoupling and the cut-off $\lambda^{c}_{\rm fsh}$, ensuring enhancement in small scale structure, provide a new handle to distinguish the parameter space of a DM model in presence of a non-standard cosmological epoch. We utilize this idea to identify novel areas in the parameter space of two simplified models: one with a scalar DM involving $s$-wave scattering by scalar quartic interactions (Model I), and the other with fermionic DM showing $p$-wave scattering by Yukawa interactions (Model II). We obtain viable region of parameter space in the WIMP DM mass range for both models, where the DM decouples partially (Model I) or fully (Model II) resulting in $\lambda_{\rm fsh}<\lambda^{c}_{\rm fsh}$, during an EMDE with temperature dependent entropy injection which yields a sub-GeV reheating temperature. This results in an amplification of sub-earth halo population with a peak around $M=0.1M_{\oplus}$ for $T_{\rm RH}=10$ MeV. In comparison, an EMDE with $\Gamma_{\phi}\sim{\rm const.}$ requires larger values of $T_{\rm RH}\sim{\rm GeV}$ to reproduce the correct relic density of the DM, which would result in formation of structures at much smaller scales($M<10^{-5}M_{\oplus}$) compared to $M=0.1M_{\oplus}$, thus providing a direct observational access to the future experiments in probing different non-standard cosmological epochs.
@@ Appendix A Perturbation equations
We provide the detailed equations governing the density and velocity perturbations of the scalar $\phi$, radiation and the DM fluids below.
@@
Here, the prime denotes differentiation with respect to the scale factor ($a$), and $\rho_{i}$’s denote the background densities. The above equations are written in the Fourier space where the $\nabla^{2}$ is replaced by $-k^{2}$.
For EMDE with time dependent entropy injection, i.e. $\Gamma_{\phi}\sim T$, we get $\delta\Gamma_{\phi}/\Gamma_{\phi}\sim\delta_{\gamma}/4$. The $\langle\sigma v\rangle_{\rm ann}$ and its derivative with respect to the temperature has to be calculated for a given DM model. For example, in the Model I of Sec. 4.1, we find
@@
Figure 10: Full numerical solution (orange) and best fit curve (blue dashed) for $\delta_{\chi}$ as a function of $k/k_{\rm RH}$ are shown at $a/a_{\rm RH}=192.5$ keeping $T_{\rm RH}=10$ MeV. For $k<k_{\rm min}$, $\delta_{\chi}$ is constant as these modes are still super-Hubble, for $k_{\rm min}<k<0.1\,k_{\rm RH}$ $\delta_{\chi}$ grows since the metric perturbation $\Phi$ has not decayed, while for $0.1\,k_{\rm RH}<k<k_{\rm RH}$, $\delta_{\chi}$ settles into logarithmic growth. Modes with $k>k_{\rm RH}$ entered the horizon during the EMDE, resulting in $\delta_{\chi}$ displaying power-law momentum dependence. The best-fit curve closely follows the full numerical solution for $k>0.1\,k_{\rm RH}$.
@@
In Fig. 10 we show that the best fit curve (in orange) matches well with the numerical solution (blue dashed curve) for $k>0.05\,k_{\rm RH}$.

### v2

@@ 4.2 Model II: p-wave scattered fermion dark matter
Figure 8: The allowed parameter region satisfying $m_{\chi}<M\leq 2m_{\chi}$ and reproducing the present day relic density of dark matter for a fixed value of $T_{\rm RH}=100$ MeV. In the dark-shaded region, dark matter kinetically decouples during the EMDE, while it decouples during radiation-dominated epoch in the light-shaded region. Enhancement in the sub-earth halo population can not be observed in the region on the left side of the cyan dashed curve, even though DM remains fully kinetically decoupled during EMDE. The isocontours of the free-streaming horizon are shown by white lines.
@@
  • The DM shows full kinetic decoupling during the EMDE in the dark-coloured area, while the relic is satisfied in the entire parameter space within the specified boundaries.
  • The isocontours of the free-streaming horizon are projected as white curves on the allowed parameter space. Due to the full decoupling of the DM, the value of $\lambda_{\rm fsh}$ is approximately two orders of magnitude smaller compared to a similarly situated benchmark point in Fig. 7, where the DM decouples only partially.
  • The region left to the cyan dot-dashed line does not exhibit a boost in the population of sub-earth halos even if DM becomes fully decoupled during the EMDE. This is because the free-streaming horizon of the DM in that region exceeds the cut-off for the corresponding $T_{\rm RH}$.
@@
We illustrate the differential halo mass function at several redshift values for both Model I and II in Fig. 9, keeping $T_{\rm RH}=10$ MeV and $m_{\chi}=95$ GeV for both cases. The relevant couplings are chosen, as in Fig. 3, to satisfy the present-day relic density. For this benchmark point, the DM decouples around $T=2.2$ GeV for both Models. However, the free-streaming horizon of the DM is $\lambda_{\rm fsh}=0.264$ pc for Model I and $0.019$ pc for Model II, respectively. The smaller $\lambda_{\rm fsh}$ in Model II is attributed to the extra cooling due to full kinetic decoupling before the end of reheating. The additional cooling in Model II is further confirmed by the DM temperature at $a_{\rm RH}$, with $T_{\chi}(a_{\rm RH})$ being $55.5$ KeV for Model I and $12$ eV for Model II. As shown in Fig. 9, the differential halo mass function peaks around $M=0.1\,M_{\oplus}$ for both models, while the absolute value is higher in Model II due to the reduced $\lambda_{\rm fsh}$.
Figure 9: Differential halo mass fraction is demonstrated for both the models at different redshift values. For the benchmark points as mentioned in the caption of Fig. 3, the free-streaming horizon of DM turns out to be $0.264$ $pc$ and $0.019$ $pc$ for Model I and Model II, respectively. The differential halo mass function peaks around $M=0.1\,M_{\oplus}$.
@@ 5 Conclusions
We have comprehensively studied the chemical and kinetic decoupling of thermal dark matter in the presence of a pre-BBN early matter-dominated era. Specifically, we explore a scenario where a long-lived scalar field dominated the energy density of the universe, eventually transferring energy to the radiation through a temperature-dependent dissipation rate ($\Gamma_{\phi}\sim T$). We demonstrate that the kinetic decoupling of the DM during the EMDE is partial if the elastic scattering of the DM with the bath particles is dominated by $s$-wave partial wave amplitudes, while the DM fully decouples during the EMDE if it undergoes $p$-wave scattering. This is in contrast to an EMDE with a constant rate of energy dissipation ($\Gamma_{\phi}\sim{\rm const.}$), where $s$-wave scatterings are inefficient to kinetically decouple the DM, while $p$-wave scatterings can partially decouple it from the thermal bath.
@@
An early kinetic decoupling reduces the free-streaming horizon of the DM ($\lambda_{\rm fsh}$) due to additional cooling of the DM during the EMDE. In addition, the presence of an EMDE also enhances the matter perturbations for scales entering the horizon during this epoch. The population of the DM halos with masses smaller than $M_{\rm RH}$ receive a boost due to the combined effect of reduced free-streaming horizon and enhanced matter perturbations. The boost in the number of small-scale DM halos may in turn lead to an increase in the gamma-ray annihilation signals from the DM halos, offering an interesting probe for the early matter-dominated cosmology. Gravitational methods, such as pulsar timing arrays, microlensing, and detection of primordial gravitational waves provide alternative prospects to probe the physics of the early matter-dominated universe.
@@
However, kinetic decoupling of the DM during the EMDE does not guarantee an enhanced structure growth. If the annihilation of the DM which determines its relic and the elastic scattering with bath particles are governed by the same interaction strength, the free-streaming horizon reduces in the EMDE. On the other hand, in the presence of multiple interactions which decorrelate the mechanisms for chemical and kinetic decoupling, the free-streaming horizon may not be reduced in the presence of the EMDE unless the kinetic decoupling temperature is greater than a threshold value compared to the reheating temperature, see Fig. 1.
@@
Moreover, we have shown that the formation of DM halos receives a boost iff the free-streaming horizon is smaller than a certain cut-off ($\lambda^{c}_{\rm fsh}$). We have quantified $\lambda^{c}_{\rm fsh}$, which depends on the reheating temperature, by comparing $\sigma(M)$ in the presence of EMDE and in the purely radiation-dominated universe with $\lambda_{\rm fsh}=0$.
@@
As mentioned above, if the chemical and kinetic decoupling are correlated by the same interaction strengths, the requirement of kinetic decoupling and the cut-off $\lambda^{c}_{\rm fsh}$, ensuring enhancement in small-scale structure, provide a new handle to distinguish the parameter space of a DM model in presence of a non-standard cosmological epoch. We utilize this idea to identify novel areas in the parameter space of two simplified models: one with a scalar DM involving $s$-wave scattering by scalar quartic interactions (Model I), and the other with fermionic DM showing $p$-wave scattering by Yukawa interactions (Model II). We obtain the viable region of parameter space in the WIMP DM mass range for both models, where the DM decouples partially (Model I) or fully (Model II) resulting in $\lambda_{\rm fsh}<\lambda^{c}_{\rm fsh}$, during an EMDE with temperature-dependent entropy injection which yields a sub-GeV reheating temperature. This results in an amplification of the sub-earth halo population with a peak around $M=0.1M_{\oplus}$ for $T_{\rm RH}=10$ MeV. In comparison, an EMDE with $\Gamma_{\phi}\sim{\rm const.}$ requires larger values of $T_{\rm RH}\sim{\rm GeV}$ to reproduce the correct relic density of the DM, which would result in the formation of structures at much smaller scales($M<10^{-5}M_{\oplus}$) compared to $M=0.1M_{\oplus}$, thus providing a direct observational access to the future experiments in probing different non-standard cosmological epochs.
@@ Appendix A Perturbation equations
We provide the detailed equations governing the density and velocity perturbations of the scalar $\phi$, radiation, and the DM fluids below.
@@
Here, the prime denotes differentiation with respect to the scale factor ($a$), and $\rho_{i}$’s denotes the background densities. The above equations are written in the Fourier space where the $\nabla^{2}$ is replaced by $-k^{2}$.
For EMDE with temperature-dependent entropy injection, i.e., $\Gamma_{\phi}\sim T$, we get $\delta\Gamma_{\phi}/\Gamma_{\phi}\sim\delta_{\gamma}/4$. The $\langle\sigma v\rangle_{\rm ann}$ and its derivative with respect to the temperature has to be calculated for a given DM model. For example, in the Model I of Sec. 4.1, we find
@@
Figure 10: Full numerical solution (orange) and best-fit curve (blue dashed) for $\delta_{\chi}$ as a function of $k/k_{\rm RH}$ are shown at $a/a_{\rm RH}=192.5$ keeping $T_{\rm RH}=10$ MeV. For $k<k_{\rm min}$, $\delta_{\chi}$ is constant as these modes are still super-Hubble, for $k_{\rm min}<k<0.1\,k_{\rm RH}$ $\delta_{\chi}$ grows since the metric perturbation $\Phi$ has not decayed, while for $0.1\,k_{\rm RH}<k<k_{\rm RH}$, $\delta_{\chi}$ settles into logarithmic growth. Modes with $k>k_{\rm RH}$ entered the horizon during the EMDE, resulting in $\delta_{\chi}$ displaying power-law momentum dependence. The best-fit curve closely follows the full numerical solution for $k>0.1\,k_{\rm RH}$.
@@
In Fig. 10 we show that the best-fit curve (in orange) matches well with the numerical solution (blue dashed curve) for $k>0.05\,k_{\rm RH}$.